If the initial angular speed of the fan blades is 3.0 rad / sec, what is the final angular speed of the fan blades in rad/sec? 


1
Expert's answer
2022-02-02T12:48:20-0500
"\omega_f=\omega_i+\alpha t=3+\alpha t,"

where "\alpha" and t are the acceleration and time.